OpenStudy (anonymous):

if costheta+sintheta=(2)^1/2 costheta, then (costheta-sintheta)=? a.(2)^1/2 costheta b..(2)^1/2 sintheta c..2 costheta d.2 sintheta

OpenStudy (anonymous):

Have you tried to solve this question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Do you know basic trigonometry?

OpenStudy (anonymous):

explaim me this question

OpenStudy (anonymous):

In this question you have to find value of (costheta - sintheta) from given value of (sintheta + costheta).

OpenStudy (anonymous):

explin mw in written

OpenStudy (anonymous):

what is mw?

OpenStudy (anonymous):

explain me in written

OpenStudy (anonymous):

For this question you first need to square costheta + sintheta=(2)^1/2 costheta Can you tell me what it will be?

OpenStudy (anonymous):

1+2sinthetacostheta=2costheta^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now take 2sinthetacostheta to the right hand side.

OpenStudy (anonymous):

Can you tell me what the equation will be now?

OpenStudy (anonymous):

1=2costheta(costheta-sintheta)

OpenStudy (anonymous):

Now you have got (costheta-sintheta) in your equation. Isolate it.

OpenStudy (anonymous):

please explain next in written

OpenStudy (anonymous):

(costheta-sintheta) = 1/2costheta